# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ C.1 Local Aggregation
We define $g^{(0)}_{j}=\frac{1}{m}\sum_{i=1}^{B}(p_{ki}^{(0)}-y_{ki}^{(0)})\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})$, $g^{(1)}_{j}=\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}^{(1)}-y_{ki}^{(1)})\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$ and $\tilde{g}_{j}=\eta(g^{(0)}_{j}+g^{(1)}_{j})$, then the target vector $\tilde{g}=\sum_{j=1}^{m}\tilde{g}_{j}$. We also define $\hat{P}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{P}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{{\bm{T}}}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{{\bm{T}}}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{P}=\eta(\hat{P}_{0}+\hat{P}_{1})$ and $\hat{{\bm{T}}}=\eta(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1})$. Let $P=\operatorname{\mathbb{E}}\hat{P}_{0}$ and ${\bm{T}}=\operatorname{\mathbb{E}}\hat{{\bm{T}}}_{0}$. We will bound $\|\hat{P}_{0}+\hat{P}_{1}-2P\|$ and $\|(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-2{\bm{T}})(V,V,V)\|$.
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Eq. (10). Now we only need to bound $\|\hat{P}_{1}-\hat{P}_{0}\|$. By Lemma C.4 and Lemma C.8,
$\|\hat{P}_{1}-$ $\hat{P}_{0}\|=\left\|\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}\left[(p_{1i}-y_{1i})+(p_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right\|$
$\leq\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\quad\quad+\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\leq\tilde{O}(\frac{B^{2}Kd}{m})+\tilde{O}(\frac{B^{2}d}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(K^{2}B^{2}d)$, $\|\hat{P}_{1}-\hat{P}_{0}\|\leq O(B\sqrt{\frac{d}{m}})$. Thus,
$\|\hat{P}-2\eta P\|=\eta\|\hat{P}_{0}+\hat{P}_{1}-2P\|\leq\eta\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Eq. (12). Now we only need to bound $\|\hat{{\bm{T}}}(V,V,V)_{1}-\hat{{\bm{T}}}_{0}(V,V,V)\|$. By Lemma C.5 and Lemma C.9,
@@
$=$ $\left\|\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}\left[(p_{1i}-y_{1i})+(p_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|$
$\quad\quad+\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\tilde{O}(\frac{B^{7/2}K}{m})+\tilde{O}(\frac{B^{7/2}}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(K^{2}B^{2}d)$, $\|\hat{{\bm{T}}}_{1}-\hat{{\bm{T}}}_{0}\|\leq O(\frac{B^{5/2}}{m})$. Thus,
@@
By Lemma 3.1, we have $P=\sum_{i=1}^{B}2(\frac{2}{K}-y_{1i}-y_{2i})\nu{\bm{x}}_{i}^{\otimes 2}$ and ${\bm{T}}=\sum_{i=1}^{B}2(\frac{2}{K}-y_{1i}-y_{2i})\lambda{\bm{x}}_{i}^{\otimes 3}$. Since $\frac{1}{K}\leq|r_{i}|\leq 1$, we have that $\eta P$’s smallest component $\nu_{\mathrm{min}}\geq\frac{2\eta|\nu|}{K}$, $\eta{\bm{T}}$’s smallest component $\lambda_{\mathrm{min}}\geq\frac{2\eta|\lambda|}{K}$ and $\kappa\leq K$. Then by Theorem A.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}_{i}}\|\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
We define $g^{(0)}_{j}=\frac{1}{m}\sum_{i=1}^{B}(p_{ki}^{(0)}-y_{ki}^{(0)})\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})$, $g^{(1)}_{j}=\frac{\eta}{m}\sum_{i=B+1}^{N}(p_{ki}^{(1)}-y_{ki}^{(1)})\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$, $\tilde{g}^{(1)}_{j}=\frac{1}{m}\sum_{i=B+1}^{N}(p_{ki}^{(0)}-y_{ki}^{(0)})\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})$ and $\tilde{g}_{j}=\eta(g^{(0)}_{j}+g^{(1)}_{j})$, then the target vector $\tilde{g}=\sum_{j=1}^{m}\tilde{g}_{j}$. We also define $\hat{P}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{P}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\tilde{P}_{1}=\sum_{j=1}^{m}\tilde{g}^{(1)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{{\bm{T}}}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{{\bm{T}}}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\tilde{{\bm{T}}}_{1}=\sum_{j=1}^{m}\tilde{g}^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{P}=\eta(\hat{P}_{0}+\hat{P}_{1})$ and $\hat{{\bm{T}}}=\eta(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1})$. Let $P=\operatorname{\mathbb{E}}[\hat{P}_{0}+\tilde{P}_{1}]$ and ${\bm{T}}=\operatorname{\mathbb{E}}[\hat{{\bm{T}}}_{0}+\tilde{{\bm{T}}}_{1}]$. We will bound $\|\hat{P}_{0}+\hat{P}_{1}-P\|$ and $\|(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-{\bm{T}})(V,V,V)\|$.
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Eq. (10). Now we only need to bound $\|\hat{P}_{1}-\tilde{P}_{1}\|$. By Lemma C.4 and Lemma C.8,
$\|\hat{P}_{1}-$ $\tilde{P}_{1}\|=\left\|\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}\left[(p_{1i}-y_{1i})+(p_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right\|$
$\leq\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\quad\quad+\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\leq\tilde{O}(\frac{(N-B)^{2}Kd}{m})+\tilde{O}(\frac{(N-B)^{2}d}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(K^{2}B^{2}d)$, $\|\hat{P}_{1}-\hat{P}_{0}\|\leq O(N\sqrt{\frac{d}{m}})$. Thus,
$\|\hat{P}-\eta P\|=\eta\|\hat{P}_{0}+\hat{P}_{1}-P\|\leq\eta\tilde{O}(N\sqrt{\frac{d}{m}})$
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Eq. (12). Now we only need to bound $\|\hat{{\bm{T}}}(V,V,V)_{1}-\hat{{\bm{T}}}_{0}(V,V,V)\|$. By Lemma C.5 and Lemma C.9,
@@
$=$ $\left\|\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}\left[(p_{1i}-y_{1i})+(p_{2i}-y_{2i})\right]\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\!\!\sum_{i=B+1}^{N}\!\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\!\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|$
$\quad\quad+\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(p_{1i}^{(1)}+p_{2i}^{(1)}-p_{1i}-p_{2i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\tilde{O}(\frac{(N-B)^{7/2}K}{m})+\tilde{O}(\frac{(N-B)^{7/2}}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(K^{2}B^{2}d)$, $\|\hat{{\bm{T}}}_{1}-\hat{{\bm{T}}}_{0}\|\leq O(\frac{N^{5/2}}{m})$. Thus,
$\|\hat{{\bm{T}}}-2\eta{\bm{T}}\|=\eta\|\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-2{\bm{T}}\|\leq\eta\tilde{O}(\frac{N^{5/2}}{m})$
@@
By Lemma 3.1, we have $P=\sum_{i=1}^{N}(\frac{2}{K}-y_{1i}-y_{2i})\nu{\bm{x}}_{i}^{\otimes 2}$ and ${\bm{T}}=\sum_{i=1}^{N}(\frac{2}{K}-y_{1i}-y_{2i})\lambda{\bm{x}}_{i}^{\otimes 3}$. Since $\frac{1}{K}\leq|r_{i}|\leq 1$, we have that $\eta P$’s smallest component $\nu_{\mathrm{min}}\geq\frac{\eta|\nu|}{K}$, $\eta{\bm{T}}$’s smallest component $\lambda_{\mathrm{min}}\geq\frac{\eta|\lambda|}{K}$ and $\kappa\leq K$. Then by Theorem A.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}_{i}}\|\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(N\sqrt{\frac{d}{m}})$

### v2

@@ C.1 Local Aggregation
We define $g^{(0)}_{j}=\frac{1}{m}\sum_{i=1}^{B}r_{i}^{(0)}\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})$, $g^{(1)}_{j}=\frac{1}{m}\sum_{i=1}^{B}r_{i}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$ and $\tilde{g}_{j}=g^{(0)}_{j}+g^{(1)}_{j}$, then the target vector $\tilde{g}=\sum_{j=1}^{m}\tilde{g}_{j}$. We also define $\hat{P}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{P}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{{\bm{T}}}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{{\bm{T}}}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{P}=\hat{P}_{0}+\hat{P}_{1}$ and $\hat{{\bm{T}}}=\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}$. Let
$P=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{B}r_{i}^{*}\sigma^{\prime\prime}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}{\bm{x}}_{i}^{\top}\right]$
and
${\bm{T}}=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{B}r_{i}^{*}\sigma^{(3)}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}^{\otimes 3}\right].$
We will bound $\|\hat{P}_{0}+\hat{P}_{1}-2P\|$ and $\|(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-2{\bm{T}})(V,V,V)\|$.
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Lemma B.4. Now we only need to bound $\|\hat{P}_{1}-\hat{P}_{0}\|$. By Lemma C.7 and Lemma C.9,
$\|\hat{P}_{1}-$ $\hat{P}_{0}\|=\left\|\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}r_{i}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)-\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right\|$
$\leq\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\quad\quad+\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\leq\tilde{O}(\frac{B^{2}d}{m})+\tilde{O}(\frac{B^{2}d}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(B^{2}d)$, $\|\hat{P}_{1}-\hat{P}_{0}\|\leq O(B\sqrt{\frac{d}{m}})$. Thus,
$\|\hat{P}-2P\|=\|\hat{P}_{0}+\hat{P}_{1}-2P\|\leq\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Lemma B.5. Now we only need to bound $\|\hat{{\bm{T}}}(V,V,V)_{1}-\hat{{\bm{T}}}_{0}(V,V,V)\|$. By Lemma C.8 and Lemma C.10,
@@
$=$ $\left\|\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}r_{i}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=1}^{B}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|$
$\quad\quad+\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\tilde{O}(\frac{B^{7/2}}{m})+\tilde{O}(\frac{B^{7/2}}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(B^{2}d)$, $\|\hat{{\bm{T}}}_{1}-\hat{{\bm{T}}}_{0}\|\leq O(\frac{B^{5/2}}{m})$. Thus,
@@
We have that $P$’s smallest component $\nu_{\mathrm{min}}\geq|\nu|$, ${\bm{T}}$’s smallest component $\lambda_{\mathrm{min}}\geq|\nu|$ and $\kappa=\frac{\max|r_{i}^{*}|}{\min|r_{i}^{*}|}=1$. Since $0<|y_{i}|\leq 1$, then $|r_{i}^{*}|$ is lower bounded. By Theorem B.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}_{i}}\|\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
We define $g^{(0)}_{j}=\frac{1}{m}\sum_{i=1}^{B}r_{i}^{(0)}\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})$, $g^{(1)}_{j}=\frac{1}{m}\sum_{i=B+1}^{N}r_{i}^{(1)}sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$, $\tilde{g}^{(1)}_{j}=\frac{1}{m}\sum_{i=B+1}^{N}r_{i}^{(0)}sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})$ and $\tilde{g}_{j}=g^{(0)}_{j}+g^{(1)}_{j}$, then the target vector $\tilde{g}=\sum_{j=1}^{m}\tilde{g}_{j}$. We also define $\hat{P}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{P}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\tilde{P}_{1}=\sum_{j=1}^{m}\tilde{g}_{j}^{(1)}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, $\hat{{\bm{T}}}_{0}=\sum_{j=1}^{m}g^{(0)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{{\bm{T}}}_{1}=\sum_{j=1}^{m}g^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\tilde{{\bm{T}}}_{1}=\sum_{j=1}^{m}\tilde{g}^{(1)}_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$, $\hat{P}=\hat{P}_{0}+\hat{P}_{1}$ and $\hat{{\bm{T}}}=\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}$. Let
$P=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{N}r_{i}^{*}\sigma^{\prime\prime}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}{\bm{x}}_{i}^{\top}\right]$
and
${\bm{T}}=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{N}r_{i}^{*}\sigma^{(3)}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}^{\otimes 3}\right].$
We will bound $\|\hat{P}_{0}+\hat{P}_{1}-2P\|$ and $\|(\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-2{\bm{T}})(V,V,V)\|$.
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Lemma B.4. Now we only need to bound $\|\hat{P}_{1}-\tilde{P}_{1}\|$. By Lemma C.7 and Lemma C.9,
$\|\hat{P}_{1}-$ $\tilde{P}_{1}\|=\left\|\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}r_{i}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right.$
$\quad\quad\quad\quad\left.-\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I\right)\right\|$
$\leq\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\quad\quad+\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|$
$\leq\tilde{O}(\frac{(N-B)^{2}d}{m})+\tilde{O}(\frac{(N-B)^{2}d}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(B^{2}d)$, $\|\hat{P}_{1}-\hat{P}_{0}\|\leq O(N\sqrt{\frac{d}{m}})$. Thus,
$\|\hat{P}-P\|=\|\hat{P}_{0}+\hat{P}_{1}-P\|\leq\tilde{O}(N\sqrt{\frac{d}{m}})$
@@
with probability $1-\frac{\delta}{2}$, where the last inequality is by Lemma B.5. Now we only need to bound $\|\hat{{\bm{T}}}(V,V,V)_{1}-\hat{{\bm{T}}}_{0}(V,V,V)\|$. By Lemma C.8 and Lemma C.10,
@@
$=$ $\left\|\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}r_{i}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right.$
$\quad\quad-\left.\frac{1}{m}\sum_{i=B+1}^{N}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\!\!\sum_{i=B+1}^{N}\!\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\!\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)(V,V,V)\right\|$
$\quad\quad+\sum_{i=B+1}^{N}\left\|\frac{1}{m}\sum_{j=1}^{m}\left(r_{i}^{(1)}-r_{i}\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\left({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I\right)(V,V,V)\right\|$
$\leq$ $\tilde{O}(\frac{(N-B)^{7/2}}{m})+\tilde{O}(\frac{(N-B)^{7/2}}{m})$
with probability $1-\frac{\delta}{2}$. Since $m\geq\tilde{\Omega}(B^{2}d)$, $\|\hat{{\bm{T}}}_{1}-\hat{{\bm{T}}}_{0}\|\leq O(\frac{N^{5/2}}{m})$. Thus,
$\|\hat{{\bm{T}}}-2{\bm{T}}\|=\|\hat{{\bm{T}}}_{0}+\hat{{\bm{T}}}_{1}-2{\bm{T}}\|\leq\tilde{O}(\frac{N^{5/2}}{m})$
@@
We have that $P$’s smallest component $\nu_{\mathrm{min}}\geq|\nu|$, ${\bm{T}}$’s smallest component $\lambda_{\mathrm{min}}\geq|\nu|$ and $\kappa=\frac{\max|r_{i}^{*}|}{\min|r_{i}^{*}|}=1$. Since $0<|y_{i}|\leq 1$, then $|r_{i}^{*}|$ is lower bounded. By Theorem B.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}_{i}}\|\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(N\sqrt{\frac{d}{m}})$